OpenStudy (anonymous):

what is the square root of 99? i had it , but went and confused myself .. help ??

OpenStudy (kidrah69):

Hi there! Welcome to openstudy!! :) Are there options? (Answer choices)

OpenStudy (anonymous):

No , i have to put these numbers in order from least to greatest .. square root of 99 , 3.14 squared and 9.88 repeating

OpenStudy (kidrah69):

the sqrt of 99 is 9.95 9.94987437107 <---this is the answer the calculator gave i rounded the answer to the nearest tenth Now you can put the numbers in order hopefully :)

OpenStudy (anonymous):

thanks :)

OpenStudy (kidrah69):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

hero (hero):

\(\sqrt{99} \approx \sqrt{100}\)

OpenStudy (mathstudent55):

Use a calculator, and write every number as a decimal. Then compare them.

OpenStudy (kidrah69):

Its better to put it in decimal for since the numbers given are already in the form of decimals

OpenStudy (kidrah69):

except the sqrt of 99

OpenStudy (mathstudent55):

\(\sqrt{99} = 9.94987437...\)

OpenStudy (triciaal):

@kidrah69 to the nearest tenth means to 1 decimal place to the nearest hundredth means to 2 decimal places